OpenStudy (anonymous):

a polynomial function p of degree 3 contains the points (0,3) (1,1) (2,9) (3,33). find an equation for p.

OpenStudy (amistre64):

if they actually do fit on a degree 3, then you only need 3 points to construct it with

OpenStudy (amistre64):

make 3 equations with 3 unknowns and solve with your favorite method

OpenStudy (amistre64):

\[\begin{array}\ x^3a+x^2b+xd+c=y\\x^3a+x^2b+xd+c=y\\x^3a+x^2b+xd+c=y\\x^3a+x^2b+xd+c=y\\\end{array}\] hmm, looks like i miscounted to begin with :) for a degree 3 we need 4 points ...

OpenStudy (amistre64):

plug in the values for each point into a separate row and then the fastest way would be to row reduce an augmented matrix form it

OpenStudy (anonymous):

augmented matrix?

OpenStudy (amistre64):

yes, which is a suitable way to obtain the missing values; but it is math level dependant ....

OpenStudy (anonymous):

im not at that level yet, how do i start?

OpenStudy (amistre64):

well, take the generic set up ive provided and fill in the point values; one row for each point. this will give you a set of 4 equation to play with

OpenStudy (anonymous):

what do i do once i have the equations?

OpenStudy (amistre64):

substitutions, or eliminations, or any other method of solving a system of equations that you ahve come across and are confident in using .... there is no single way to approach this.

OpenStudy (amistre64):

that (0,y) point will provide you a useful way to knock this down to a system of 3 equations right off the bat

OpenStudy (amistre64):

0a + 0b + 0c + d = 3 ; therefore d=3 can be reused on the others and we are down to 3 equations to solve for